# Quantum-limited superresolution of two incoherent point sources with unknown photon numbers

### v1

@@ IV Superresolution for two arbitrary point sources
In practical applications, as the two photon sources are incoherent, the point-spread functions of the two sources can differ from each other in general. Such a difference in the point-spread functions reduces the overlap between the two optical sources, so an intuitive idea is that it may increase the distinguishability of the two sources which could be helpful to the estimation precision of quantum superresolution for the centroid separation of the two sources. In the following, we generalize the result of the previous section and study the condition of quantum superresolution for two different optical point sources.
@@
The estimation precision for the centroid separation can still be obtained by the scalar quantum Cramér-Rao bound (18). The quantum Fisher information matrix with respect to the unknown parameters $\bar{X},d,\epsilon$ turns out to be
@@
Considering the centroid separation $d$ is small, we expand the numerator and the denominator of $\mathcal{H}_{d}$ to the fourth order of $d$ respectively ,
@@
Table 1: Summary of the ratios $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ between the estimation precision $\mathcal{H}_{d}$ of the centroid separation $d$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (37) in different regimes of the parameters. The difference between the point-spread functions and that between the photon numbers are characterized by Eq. (43) and we denote $1-\langle\psi_{1}|\psi_{2}\rangle=b\text{$\widetilde{d}^{h}$}$, $\frac{1}{2}-\xi=a\widetilde{d}^{f}$, $\chi=c\widetilde{d}^{e}$, $\epsilon=y\widetilde{d}^{s}$, where $y,a,b,c$ are real coefficients, the exponents $s,f,h,e$ are nonnegative integers and $h,\,f,\,e$ satisfy the constraint relation (44). The blue region represents the cases that attain the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$, the red region represents the cases that the estimation precision vanishes, and the yellow region represents the cases that the quantum superresolution can still work but with a lower precision. The estimation precision of the centroid separation given by the original superresolution scheme in [20] corresponds to $h\geq 3$, $s\geq 2$, and the resurgence of Rayleigh’s curse [85] corresponds to $h\geq 3,\,s=0$. The total photon number $N_{{\rm tot}}$ is sufficiently large, i.e., $N_{{\rm tot}}\gg 1$. The results are evaluated in the limit $d\rightarrow 0$. As the two point-spread functions approach to be the same in this limit, the subscripts of $\left\langle P^{4}\right\rangle{}_{12}$ and $\left\langle P^{2}\right\rangle{}_{12}$ are dropped in the table.
@@
$\ h\geq 5\\ \\ \end{gathered}$
$\ \frac{{\rm Var}(\hat{P}^{2})}{4y^{2}\langle\hat{P}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}\\ \\ \end{gathered}$
$\ h=4\\ \\ \end{gathered}$
$\ \frac{8b\langle\hat{P}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}{4(2b+y^{2})\langle\hat{P}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}\\ \\ \end{gathered}$ $\begin{gathered}0\\ \\ \\ \end{gathered}$
@@
$\ h=2\\ \\ \end{gathered}$
$\ 1\\ \\ \end{gathered}$ $\begin{gathered}\frac{2b\left(1-y^{2}\right)}{2b+y^{2}}\end{gathered}$
$\ h=1\\ \\ \end{gathered}$
$\begin{gathered}1-y^{2}\end{gathered}$
$\ h=0\\ \\ \end{gathered}$
$\begin{gathered}1-c^{2}\end{gathered}$ $\begin{gathered}\frac{\left(1-c^{2}\right)\left(1-y^{2}\right)}{1+cy}\end{gathered}$
Similarly, it is shown in Appendix D that the coefficients $A_{2}$ and $B_{2}$ of the quadratic terms in $\mathcal{H}_{d}$ may also become small but nonzero when the two point-distribution functions differ slightly from each other, so if they are of order $O(d^{2})$, they will also compete with the fourth-order terms in determining $\mathcal{H}_{d}$. Moreover, the difference between the photon numbers of the two optical sources are also involved in $\mathcal{H}_{d}$ which may compete with the higher-order terms in $\mathcal{H}_{d}$ if the difference in the photon numbers of the two point sources is small. Combining all these together, the estimation precision of the centroid separation and thus the validity of superresolution are highly dependent on the competition between the difference in the two point-spread functions, the difference in the photon numbers of the two point sources and the magnitude of the centroid separation.
@@
But note that the photon numbers and the closeness between the point-spread functions are dimensionless while the separation $d$ is not, so for the sake of convenience to make these quantities comparable and characterize the magnitudes of these quantities, we rescale the centroid separation $d$ to make it dimensionless. We define the following dimensionless $\widetilde{d}$,
@@
where $V=1/\sqrt{\langle\hat{P}^{2}\rangle}$. The subscript is dropped from $\langle\hat{P}^{2}\rangle$ in the definition of $V$, as the following results of the estimation precision $\mathcal{H}_{d}$ are evaluated in the limit $d\rightarrow 0$ which makes $\langle\hat{P}^{2}\rangle_{1}=\langle\hat{P}^{2}\rangle_{2}=\langle\hat{P}^{2}\rangle_{12}$. In order to represent the magnitude of these quantities which are all considered to be small, we denote them in the orders of $\widetilde{d}$,
@@
A detailed proof of this relation can be found in Appendix E.2. This constraint will be taken into account in the following computation of $\mathcal{H}_{d}$. When the exponents $h,f,e,s$ vary, the estimation precision $\mathcal{H}_{d}$ of the centroid separation $d$ will be different. The estimation precision $\mathcal{H}_{d}$ can be computed by Eq. (38) for different ranges of these exponents.
The results of the estimation precision $\mathcal{H}_{d}$ are summarized in Table 1 in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (37), which covers all the possible ranges of the difference between the two point-spread functions and the difference between the two photon numbers and rules out the cases that the exponents do not satisfy the constraint relation (44). It can be seen that the superresolution protocol can actually work for a wide range of point-spread functions with various magnitudes of the photon numbers, and the resurgence of Rayleigh’s curse occur only when the order of $1-\langle\psi_{1}|\psi_{2}\rangle$ is higher than or equal to $O(\widetilde{d}^{3})$ and $\epsilon=O(1)$, which is the cases considered in Ref. [85]. The case $h\geq 3$ with $s\geq 2$ is the regime that was considered by the original quantum superresolution scheme [20]. The case $h\geq 5$ with $s=1$ is the regime where the effective condition of quantum superresolution for two identical point-spread functions was obtained in Sec. III. For the other cases, the quantum superresolution scheme can still work, but the estimation precision of the centroid separation may deviate from that of the original superresolution scheme [20].
It needs to be stressed again that the orders of the difference between the point-spread functions and the difference between the photon numbers of the two point sources in terms of the (dimensionless) centroid separation $d$ do not mean that they change with $d$ but rather impose a limitation on the magnitudes of those differences to realize the superresolution.
@@ V Example: Point sources of Gaussian distributions
Figure 1: Illustration of two Gaussian point-spread functions $\psi_{1}(x)$ and $\psi_{2}(x)$ with centroid positions $X_{1}$, $X_{2}$ and widths $\sigma_{1}$, $\sigma_{2}$ respectively. The purpose of the quantum superresolution scheme is to estimate the centroid separation $d=X_{2}-X_{1}$.
@@ V.1 Gaussian point-spread functions with the same width
For two Gaussian point-spread functions with the same width $\sigma$, the estimation precision of the centroid separation is given by Eq. (29), where the coefficients can be simplified by the Gaussian point-spread function and the estimation precision of $d$ becomes
@@
It can be seen that the ratio between $\mathcal{H}_{d}$ and $\mathcal{H}_{d}^{({\rm opt})}$ is determined by two parameters only, $\epsilon$, i.e., the ratio between the difference and the total of the two photon numbers, and $d/\sigma$, i.e., the ratio between the centroid difference and the width of the two point sources.
Figure 2: Plot of the ratio of the estimation precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ with respect to the ratio $\epsilon$ between the difference and the total of the photon numbers for two Gaussian point-spread functions given different ratios $d/\sigma$. $d$ is the centroid separation of the two point sources, and $\sigma$ is the width of the two Gaussian point-spread functions. It can be observed from the figure that when the photon numbers of the two sources are the same, i.e., $\epsilon=0$, the estimation precision can always reach the optimum, and when $\epsilon>0$, the estimation precision decreases. The ratio $d/\sigma$ influences how fast the estimation precision decays to zero, i.e., the quantum superresolution transits to the Rayleigh’s curse, when the photon number difference increases. The large the ratio $d/\sigma$ is, the slower the precision $\mathcal{H}_{d}$ decays to zero. The exact effect of the ratio $d/\sigma$ on the estimation precision $\mathcal{H}_{d}$ is determined by Eq. (102) in Appendix C.2.
Hence, if the difference between the two photon numbers keeps finite compared to the total photon number, i.e., $\epsilon=O(1)$, the estimation precision $\mathcal{H}_{d}$ vanishes as $d$ goes to 00, which is the resurgence of the Rayleigh’s curse. But if the order of the ratio $\epsilon$ is higher than or equal to $O(d/\sigma)$, the magnitude of the denominator of $\mathcal{H}_{d}$ becomes comparable to the numerator, so the estimation precision $\mathcal{H}_{d}$ will be independent of $d$, and one can still acquire information about the centroid separation in this case. In particular, if the order of $\epsilon$ is higher than $O(d/\sigma)$, $\mathcal{H}_{d}$ will be restored to the original estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with equal photon numbers obtained in [20]. If the order of $\epsilon$ is equal to $O(d/\sigma)$, the precision $\mathcal{H}_{d}$ does not vanish, but will be decreased by a factor (32),
@@
This explicitly shows the influence of the photon number difference in the estimation precision of quantum superresolution.
To illustrate the above result numerically, the ratio of the precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ (47) for two Gaussian point-spread functions with the same width is plotted with respect to $\epsilon$ for different $d/\sigma$ in Fig. 2. It can be seen that for a fixed centroid separation, the precision of the separation reaches the maximum when $\epsilon=0$, i.e., the photon numbers of the two point sources are equal, which is the regime of the original superresolution in [20], and decreases to zero when the photon number difference increases, which is the resurgence of Rayleigh’s curse found in [85]. It also shows that when $\epsilon$ is nonzero but sufficiently small, the precision $\mathcal{H}_{d}$ will still be significant and the superresolution can work, showing the effective range of $\mathcal{H}_{d}$ to avoid the resurgence of Rayleigh’s curse. Moreover, the figure also suggests that for the same $\epsilon$, a smaller centroid separation compared to the width of the Gaussian point-spread function leads to a lower estimation precision for $d$. This can be understood as for a smaller $d$, the magnitude of $\epsilon$ becomes larger compared to $d/\sigma$, so it is more difficult to satisfy the effective condition of the superresolution and the estimation precision of $d$ decreases as a result.
@@ V.2 Gaussian point-spread functions with arbitrary widths $\sigma_{1}$ and $\sigma_{2}$
Eq. (50) tells that the ratio between $\mathcal{H}_{d}$ and the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ is determined by the ration $\eta$ between the difference and the total of the widths of the two Gaussian point-spread functions and the ratio $\epsilon$ between the difference and the total of the two photon numbers only.
@@
It is worth mentioning that there is a trade-off relation between the ratios $\epsilon$ and $\eta$ in the estimation precision $\mathcal{H}_{d}$ (50). By optimizing $\mathcal{H}_{d}$ (50) over all $\epsilon$ and $\eta$, it can found that the estimation of separation $d$ reaches the highest precision when $\epsilon=\eta$, and the precision $\mathcal{H}_{d}$ is exactly the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ (52) in this case. This implies that the optimal precision of estimating the centroid separation can even be achieved when the photon numbers of the two point sources are different, if the widths of the two Gaussian point-spread functions are also different and the condition $\epsilon=\eta$ is satisfied.
Table 2: Summary of the ratios $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ between the estimation precisions $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (52) for Gaussian point-spread functions with width difference $\eta=z\widetilde{d}^{t}$ and photon number difference $\epsilon=y\widetilde{d}^{s}$. The blue region denotes the cases that reach the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$, the red region denotes the cases that the estimation precision vanishes, and the yellow region denotes the cases that the quantum superresolution work with a lower precision. The case $t\geq 2$, $s\geq 2$, where the magnitudes of the coefficients $z$ and $y$ are $O(1)$, includes the original superresolution scheme for two Gaussian point spread functions, and the case $t\geq 2$, $s=0$ is the resurgence of Rayleigh’s curse where the estimation precision of the centroid separation vanishes. All the results are evaluated in the limit $d\rightarrow 0$.
$\ s\geq 2\\ \\ \end{gathered}$
$\ s=1\\ \\ \end{gathered}$ $\ s=0\\ \\ \end{gathered}$
$\ t>2\\ \\ \end{gathered}$
[…]
$\ t=2\\ \\ \end{gathered}$
[…]
$\ \frac{1+4z^{2}}{1+4z^{2}+2y^{2}}\\ \\ \end{gathered}$
$\ t=1\\ \\ \end{gathered}$
$\ \frac{2\left(1-y^{2}\right)z^{2}}{y^{2}+2z^{2}}\\ \\ \end{gathered}$
$\ t=0\\ \\ \end{gathered}$
$\ \frac{1}{1+z^{2}}\\ \\ \end{gathered}$ $\ \frac{1-y^{2}}{1-2yz+z^{2}}\\ \\ \end{gathered}$
Figure 3: Plot of the ratios between the estimation precision $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with respect to the ratio $\epsilon$ between the difference and the total of the photon numbers for two arbitrary Gaussian point-spread functions. (a) The estimation precision $\mathcal{H}_{d}$ for different ratios $\eta$ of the width difference to the total widths of the two sources in the limit $d\rightarrow 0$. It can be observed that the estimation precision can still reach the maximum though the photon numbers of the two point sources are different, and the highest precision is achieved when $\epsilon=\eta$. The transition to the Rayleigh’s curse occurs when the difference of photon numbers further increases beyond the point $\epsilon=\eta$, and the larger $\eta$ is, the faster the estimation precision $\mathcal{H}_{d}$ decays with $\epsilon$. The exact effect of $\eta$ on the decay of the estimation precision is determined by Eq. (50). (b) The estimation precision for different ratios of the centroid separation $d$ to the total widths of the Gaussian point-spread functions $\sigma_{{\rm tot}}$ with $\eta=0.005$. The estimation precision can always reach the maximum when the photon numbers of the two sources are equal, i.e., $\epsilon=0$, and decays to zero as $\epsilon$ increases. The ratio $d/\sigma_{{\rm tot}}$ changes the decay speed of the estimation precision. The exact effect of the ratio $d/\sigma_{{\rm tot}}$ on the estimation precision is provided in Appendix (D).
@@
It should be noted that such a result of $\mathcal{H}_{d}$ in the limit $d\rightarrow 0$ implicitly assumes that the order of the leading terms $A_{0}$ and $B_{0}$ in $\mathcal{H}_{d}$ (38) is $O(1)$, i.e., the width difference $\sigma_{2}-\sigma_{1}$ is sufficiently large compared to the centroid separation $d$. When the two Gaussian optical sources are just slightly different in the widths, i.e., $\sigma_{2}-\sigma_{1}$ is small, higher-order terms of $d$ in $\mathcal{H}_{d}$ can also come into play in determining $\mathcal{H}_{d}$ when $d\rightarrow 0$. Similar to the preceding section, the precision $\mathcal{H}_{d}$ for the separation of the two centroids will be determined by the competition between the difference in the widths of Gaussian point-spread functions and difference in the photon numbers between the two optical sources in this case.
@@
where $\widetilde{d}=\frac{d}{V}$, $V=1/\sqrt{\langle\hat{P}^{2}\rangle}=\sigma_{\text{tot}}$, $s,t$ are nonnegative integers, so that we can compare the magnitudes of the photon number difference and the width difference and obtain the estimation precision $\mathcal{H}_{d}$ of the centroid separation $d$ in different ranges of the above two ratios.

### v2

@@ IV Superresolution for two arbitrary point sources
While the point-spread functions can be identical for the point sources at different positions of the object plane if the imaging system is spatially invariant with specific conditions such as the paraxial approximation [20, 27, 103], the point-spread functions can vary with the positions of the point sources in general [104]. For example, the longitudinal shifts from the focused object plane will make the locations of two point sources deviate in the depth from each other, and thus lead to different point-spread functions due to the diffraction and geometrical blurring effects [105, 106, 107]. Such a difference in the point-spread functions may reduce the overlap between the two optical sources, so an intuitive idea is that it may increase the distinguishability of the two sources which could be helpful to the resolution of the two sources. In the following, we generalize the result of the previous section and study the condition of quantum superresolution for two optical point sources with arbitrary point-spread functions.
@@
The estimation precision for the separation can still be obtained by the scalar quantum Cramér-Rao bound (19). The quantum Fisher information matrix with respect to the unknown parameters $\bar{X},d,\epsilon$ turns out to be
@@
Considering the separation $d$ is small, we expand the numerator and the denominator of $\mathcal{H}_{d}$ to the fourth order of $d$ respectively,
@@
Similarly, it is shown in Appendix D that the coefficients $A_{2}$ and $B_{2}$ of the quadratic terms in $\mathcal{H}_{d}$ may also become small but nonzero when the two point-distribution functions differ slightly from each other, so if they are of order $O(d^{2})$, they will also compete with the fourth-order terms in determining $\mathcal{H}_{d}$. Moreover, the difference between the photon numbers of the two optical sources are also involved in $\mathcal{H}_{d}$ which may compete with the higher-order terms in $\mathcal{H}_{d}$ if the difference in the photon numbers of the two point sources is small. Combining all these together, the estimation precision of the separation $d$ and thus the validity of superresolution are highly dependent on the competition between the difference in the two point-spread functions, the difference in the photon numbers of the two point sources and the magnitude of the separation.
Table 1: Summary of the ratios $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ between the estimation precision $\mathcal{H}_{d}$ of the separation $d$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (38) in different regimes of the parameters. The difference between the point-spread functions and that between the photon numbers are characterized by Eq. (44) and we denote $1-\langle\psi_{1}|\psi_{2}\rangle=b\text{$\widetilde{d}^{h}$}$, $\frac{1}{2}-\xi=a\widetilde{d}^{f}$, $\chi=c\widetilde{d}^{e}$, $\epsilon=y\widetilde{d}^{s}$, where $y,a,b,c$ are real coefficients, and $\widetilde{d}$ is the dimensionless separation (43). The exponents $s,f,h,e$ are nonnegative integers, and $h,\,f,\,e$ satisfy the constraint relation (45). The blue region represents the cases that attain the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$, the red region represents the cases that the estimation precision vanishes, and the yellow region represents the cases that the superresolution can still work but with a lower precision. The estimation precision of the separation $d$ given by the original superresolution scheme in [20] corresponds to $h\geq 3$, $s\geq 2$, and the resurgence of Rayleigh’s curse [85] corresponds to $h\geq 3,\,s=0$. The total photon number $N_{{\rm tot}}$ is sufficiently large, i.e., $N_{{\rm tot}}\gg 1$. The results are evaluated in the limit $d\rightarrow 0$. As the two point-spread functions approach to be the same in this limit, the subscripts of $\left\langle P^{4}\right\rangle{}_{12}$ and $\left\langle P^{2}\right\rangle{}_{12}$ are dropped in the table.
@@
$\ h\geq 5\\ \\ \end{gathered}$ $\ \frac{{\rm Var}(\hat{P}^{2})}{4y^{2}\langle\hat{P}{}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}\\ \\ \end{gathered}$
$\ h=4\\ \\ \end{gathered}$ $\ \frac{8b\langle\hat{P}{}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}{4(2b+y^{2})\langle\hat{P}{}^{2}\rangle^{2}+{\rm Var}(\hat{P}^{2})}\\ \\ \end{gathered}$ $\begin{gathered}0\\ \\ \\ \end{gathered}$
@@
$\ h=2\\ \\ \end{gathered}$ $\ 1\\ \\ \end{gathered}$ $\begin{gathered}\frac{2b\left(1-y^{2}\right)}{2b+y^{2}}\end{gathered}$
$\ h=1\\ \\ \end{gathered}$ $\begin{gathered}1-y^{2}\end{gathered}$
$\ h=0\\ \\ \end{gathered}$ $\begin{gathered}1-c^{2}\end{gathered}$ $\begin{gathered}\frac{\left(1-c^{2}\right)\left(1-y^{2}\right)}{1+cy}\end{gathered}$
@@
But note that the photon numbers and the closeness between the point-spread functions are dimensionless while the separation $d$ is not, so for the sake of convenience to make these quantities comparable and characterize the magnitudes of these quantities, we rescale the separation $d$ to make it dimensionless. We define the following dimensionless separation $\widetilde{d}$,
@@
where $V=1/\sqrt{\langle\hat{P}{}^{2}\rangle}$. The subscript is dropped from $\langle\hat{P}{}^{2}\rangle$ in the definition of $V$, as the following results of the estimation precision $\mathcal{H}_{d}$ are evaluated in the limit $d\rightarrow 0$ which leads to $\psi_{1}(x)=\psi_{2}(x)$ (as the difference between the two point-spread functions is characterized in terms of the power of $\widetilde{d}$ later) and thus $\langle\hat{P}{}^{2}\rangle_{1}=\langle\hat{P}{}^{2}\rangle_{2}=\langle\hat{P}{}^{2}\rangle_{12}$. In order to represent the magnitudes of these quantities which are all considered to be small, we denote them in the orders of $\widetilde{d}$,
@@
A detailed proof of this relation can be found in Appendix E.2. This constraint will be taken into account in the following computation of $\mathcal{H}_{d}$. When the exponents $h,f,e,s$ vary, the estimation precision $\mathcal{H}_{d}$ of the separation $d$ will be different. The estimation precision $\mathcal{H}_{d}$ can be computed by Eq. (39) for different ranges of these exponents.
The results of the estimation precision $\mathcal{H}_{d}$ are summarized in Table 1 in terms of the ratio between $\mathcal{H}_{d}$ and the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ (38), which covers all the possible ranges of the difference between the two point-spread functions and the difference between the two photon numbers and rules out the cases that the exponents do not satisfy the constraint relation (45). It can be seen that the superresolution protocol can actually work for a wide range of point-spread functions with various magnitudes of the photon numbers, and the estimation precision $\mathcal{H}_{d}$ vanishes only when the order of $1-\langle\psi_{1}|\psi_{2}\rangle$ is higher than or equal to $O(\widetilde{d}^{3})$ and $\epsilon=O(1)$, which is the cases considered in Ref. [85]. The case $h\geq 3$ with $s\geq 2$ is the regime that was considered by the original superresolution scheme [20]. The case $h\geq 5$ with $s=1$ is the regime where the effective condition of quantum superresolution for two identical point-spread functions was obtained in Sec. III. For the other cases, the superresolution scheme can still work, but the estimation precision of the separation $d$ may deviate from that of the original superresolution scheme [20].
It needs to be stressed again that the orders of the difference between the point-spread functions and the difference between the photon numbers of the two point sources in terms of the (dimensionless) separation do not mean that they change with $d$ but rather impose a limitation on the magnitudes of those differences to realize the superresolution.
@@ V.1 Gaussian point-spread functions with the same width
For two Gaussian point-spread functions with the same width $\sigma$, the estimation precision of the separation is given by Eq. (30), where the coefficients can be simplified by the Gaussian point-spread function and the estimation precision of $d$ becomes
@@
It can be seen that the ratio between $\mathcal{H}_{d}$ and $\mathcal{H}_{d}^{({\rm opt})}$ is determined by two parameters only, $\epsilon$, i.e., the ratio between the difference and the total of the two photon numbers, and $d/\sigma$, i.e., the ratio between the separation and the width of the two point sources.
Figure 1: Illustration of two Gaussian point-spread functions $\psi_{1}(x)$ and $\psi_{2}(x)$ with positions $X_{1}$, $X_{2}$ and widths $\sigma_{1}$, $\sigma_{2}$ respectively. The purpose of the superresolution scheme is to estimate the separation $d=X_{2}-X_{1}$ when $d$ is small.
Hence, if the difference between the two photon numbers keeps finite compared to the total photon number, i.e., $\epsilon=O(1)$, the estimation precision $\mathcal{H}_{d}$ vanishes as $d$ goes to 00, implying the Rayleigh’s curse applies again. But if the order of the ratio $\epsilon$ is higher than or equal to $O(d/\sigma)$, the magnitude of the denominator of $\mathcal{H}_{d}$ becomes comparable to the numerator, so the estimation precision $\mathcal{H}_{d}$ will be independent of $d$, and one can still acquire information about the separation $d$ in this case. In particular, if the order of $\epsilon$ is higher than $O(d/\sigma)$, $\mathcal{H}_{d}$ will be restored to the original estimation precision $\mathcal{H}_{d}^{({\rm opt})}$ with equal photon numbers obtained in [20]. If the order of $\epsilon$ is equal to $O(d/\sigma)$, the precision $\mathcal{H}_{d}$ does not vanish, but will be decreased by a factor (33),
@@
This explicitly shows the influence of the photon number difference in the estimation precision of superresolution.
To illustrate the above result numerically, the ratio of the precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ (48) for two Gaussian point-spread functions with the same width is plotted with respect to $\epsilon$ for different $d/\sigma$ in Fig. 2. It can be seen that for a fixed separation, the precision of the separation reaches the maximum when $\epsilon=0$, i.e., the photon numbers of the two point sources are equal, which is the regime of the original superresolution in [20], and decreases to zero when the photon number difference increases, which is the reoccurrence of Rayleigh’s curse found in [85]. It also shows that when $\epsilon$ is nonzero but sufficiently small, the precision $\mathcal{H}_{d}$ will still be significant and the superresolution can work, showing the effective range of quantum superresolution which is broader than $\epsilon=0$. Moreover, the figure also suggests that for the same $\epsilon$, a smaller separation compared to the width of the Gaussian point-spread function leads to a lower estimation precision for $d$. This can be understood as for a smaller $d$, the magnitude of $\epsilon$ becomes larger compared to $d/\sigma$, so it is more difficult to satisfy the effective condition of the superresolution and the estimation precision of $d$ decreases as a result.
@@ V.2 Gaussian point-spread functions with arbitrary widths $\sigma_{1}$ and $\sigma_{2}$
Figure 2: Plot of the ratio of the estimation precision $\mathcal{H}_{d}$ to the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ with respect to the ratio $\epsilon$ for two Gaussian point-spread functions given different ratios $d/\sigma$. $d$ is the separation of the two point sources, and $\sigma$ is the width of the two Gaussian point-spread functions. It can be observed from the figure that when the photon numbers of the two sources are the same, i.e., $\epsilon=0$, the estimation precision can always reach the optimum, and when $\epsilon>0$, the estimation precision decreases. The ratio $d/\sigma$ influences how fast the precision $\mathcal{H}_{d}$ decays to zero, i.e., the speed that $\mathcal{H}_{d}$ transits from the regime of quantum superresolution to the regime of Rayleigh’s curse, when the photon number difference increases. It verifies Eq. (62) that $d/\sigma=0$ is the minimal point of $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$, so the large the ratio $d/\sigma$ is, the higher the precision $\mathcal{H}_{d}$ is and hence the slower $\mathcal{H}_{d}$ decays to zero, and vice versa.
@@
Eq. (51) tells that the ratio $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ is determined by the ration $\eta$ between the difference and the total of the widths of the two Gaussian point-spread functions and the ratio $\epsilon$ between the difference and the total of the two photon numbers only.
@@
It is worth mentioning that there is a trade-off relation between the ratios $\epsilon$ and $\eta$ in the estimation precision $\mathcal{H}_{d}$ (51). By optimizing $\mathcal{H}_{d}$ (51) over all $\epsilon$ and $\eta$, it can found that the estimation of the separation $d$ reaches the highest precision when $\epsilon=\eta$, and the precision $\mathcal{H}_{d}$ is exactly the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ (53) in this case. This implies that the optimal precision of estimating the separation $d$ can even be achieved when the photon numbers of the two point sources are different, if the widths of the two Gaussian point-spread functions are also different and the condition $\epsilon=\eta$ is satisfied.
@@
It should be noted that such a result of $\mathcal{H}_{d}$ in the limit $d\rightarrow 0$ implicitly assumes that the order of the leading terms $A_{0}$ and $B_{0}$ in $\mathcal{H}_{d}$ (39) is $O(1)$, i.e., the width difference $\sigma_{2}-\sigma_{1}$ is sufficiently large compared to the separation $d$. When the two Gaussian optical sources are just slightly different in the widths, i.e., $\sigma_{2}-\sigma_{1}$ is small, higher-order terms of $d$ in $\mathcal{H}_{d}$ can also come into play in determining $\mathcal{H}_{d}$ when $d\rightarrow 0$. Similar to the preceding section, the precision $\mathcal{H}_{d}$ for the separation $d$ will be determined by the competition between the difference in the widths of Gaussian point-spread functions and the difference in the photon numbers in this case.
@@
so that we can compare the magnitudes of the photon number difference and the width difference and obtain the estimation precision $\mathcal{H}_{d}$ of the separation $d$ in different ranges of the these parameters. For the simplicity of computation, we redefine the dimensionless separation $\widetilde{d}$(43) as
$\widetilde{d}=\frac{d}{(\sigma_{1}+\sigma_{2})/2}.$ (55)
In the limit $d\rightarrow 0$, $\sigma_{1}=\sigma_{2}$ as the difference between the two point-spread functions is quantified in terms of the power of $\widetilde{d}$ (54), and $\widetilde{d}$ can now be simplified to
$\widetilde{d}=\frac{d}{\sigma}.$ (56)
Table 2: Summary of the ratios $\mathcal{H}_{d}/\mathcal{H}_{d}^{({\rm opt})}$ (53) for Gaussian point-spread functions with width difference $\eta=z\widetilde{d}^{t}$ and photon number difference $\epsilon=y\widetilde{d}^{s}$. The blue region denotes the cases that reach the optimal estimation precision $\mathcal{H}_{d}^{({\rm opt})}$, the red region denotes the cases that the estimation precision vanishes, and the yellow region denotes the cases that the superresolution work with a lower precision. The case $t\geq 2$, $s\geq 2$, where the magnitudes of the coefficients $z$ and $y$ are $O(1)$, includes the original superresolution scheme for two Gaussian point spread functions, and the case $t\geq 2$, $s=0$ is the resurgence of Rayleigh’s curse where the estimation precision of the separation vanishes. All the results are evaluated in the limit $d\rightarrow 0$, and the dimensionless separation $\widetilde{d}$ is given by Eq. (56).
$\ s\geq 2\\ \\ \end{gathered}$ $\ s=1\\ \\ \end{gathered}$ $\ s=0\\ \\ \end{gathered}$
[…]
$\ t=1\\ \\ \end{gathered}$ $\ \frac{8\left(1-y^{2}\right)z^{2}}{y^{2}+8z^{2}}\\ \\ \end{gathered}$
$\ t=0\\ \\ \end{gathered}$ $\ \frac{1}{1+z^{2}}\\ \\ \end{gathered}$ $\ \frac{1-y^{2}}{1-2yz+z^{2}}\\ \\ \end{gathered}$